Keegan is 96 years younger than Lindsey. 7 years ago, Lindsey's age was 4 times Keegan's age. How old is Keegan now?

Answer by VFBundy(438)

Keegan's age now = x
Lindsey's age now = x + 96

"7 years ago, Lindsey's age was 4 times Keegan's age":

(x + 96) - 7 = 4(x - 7)

Solve for x:

x + 96 - 7 = 4x - 28

x + 89 = 4x - 28

3x = 117

x = 39

Keegan's age now = x = 39

Double-check:

Keegan's age now = x = 39
Lindsey's age now = x + 96 = 135

"Keegan is 96 years younger than Lindsey":

135 - 39 = 96

"7 years ago, Lindsey's age was 4 times Keegan's age":

(135 - 7) = 4(39 - 7)

128 = 4(32)

128 = 128